OpenStudy (idku):

HELP!! Hard calculus problem. (If that adds anything, my instructor classified it as one-variable optimization)

OpenStudy (idku):

First, I am to find the equation of the line with, (1) y-intercept: n (2) x-intercept: 16-n (where n=[1,15] ) The slope: \(\color{blue}{\displaystyle m=\frac{n-0}{0-(16-n)}=\frac{n}{n-16}:\quad n\in [1,15] }\) The equation of the line: \(\color{blue}{\displaystyle y=\left(\frac{n}{n-16}\right)x+n:\quad n\in [1,15] }\) So, I have my (linear) function: \(\color{blue}{\displaystyle f(x,n)=\left(\frac{n}{n-16}\right)x+n:\quad n\in [1,15] }\) And there is the graph of this function: ` https://www.desmos.com/calculator/hvvsu8mvur `

OpenStudy (idku):

Now, I have to find the shape (using calculus). My instructor noted that I don't need Gradients for this, just I need to know partial derivatives (which I am comfortable with).

OpenStudy (idku):

I mean that I need to find a function of a single variable (using calculus) to represent this shape (that I graphed on desmos).

OpenStudy (baru):

good god! i HAVE NO IDEA HOW I GOT THIS, i was fumbling with this problem for quite a while plot this function on the same graph and see y=16+ x - 8\(\sqrt{x}\)

OpenStudy (baru):

i am too sleepy to explain my reasoning, i am not sure if its right, i am very sleepy, and it may very well be a fluke. i will have a closer look again tomorrow.

OpenStudy (baru):

now i am too excited to sleep >.<

OpenStudy (baru):

if you zoom out and observe the plot, you will notice that each of the lines you plotted is tangent to some curve at exactly one point. we need to find that curve. lets call it y=f(x) since each line is tangent to the curve exactly once, we can assume 'n' to be a funciton of 'x' as in, for the curve y=f(x) we have a tangent at each x, and each x has a corresponding value of 'n' so \[\color{blue}{\displaystyle y=\left(\frac{n(x)}{n(x)-16}\right)x+n(x):\quad n\in [1,15] }\]

OpenStudy (baru):

another thing is, you took integer values for n just so that you could plot it. so lets drop that. n can take any value just like x.

OpenStudy (idku):

(Sorry was doing my linear algebra quiz just now ... yes, I have followed everything so far, and by the way, thanks a lot for replying)

OpenStudy (baru):

differentiate the above equation w.r.t 'x' \[\color{blue}{\displaystyle dy/dx=\left(\frac{n'(x)(n(x)-16) - (n'(x)n(x))}{(n(x)-16)^2}\right)x+\left(\frac{n(x)}{n(x)-16}\right)+n'(x) }\]

OpenStudy (idku):

ok, product rule ... :)

OpenStudy (baru):

try graphing the equation, and tell me is that what you mean when you say find the shape?

OpenStudy (baru):

i mean the one in my first comment

OpenStudy (idku):

Yes, I figured, you referred to graph y=16+ x - 8√x

OpenStudy (baru):

so, that is what you are looking for right?

OpenStudy (idku):

Yes, now I have to gather myself up and explain why this is so.

OpenStudy (idku):

i gtg now

OpenStudy (baru):

ok, i'll post my solution anyway. so continuing where i left off \[\color{blue}{\displaystyle dy/dx=\left(\frac{n'(x)(n(x)-16) - (n'(x)n(x))}{(n(x)-16)^2}\right)x+\left(\frac{n(x)}{n(x)-16}\right)+n'(x) }\]

OpenStudy (baru):

now this part is a bit tricky we know the curve y=f(x) has the tangent \( \color{blue}{\displaystyle y=\left(\frac{n(x)}{n(x)-16}\right)x+n (x) }\) at at any (x,y) so the slope of the tangent at any point on y=f(x) is \(\color{blue}{\displaystyle \left(\frac{n(x)}{n(x)-16}\right)}\) in other words \(\color{blue}{\displaystyle dy/dx= \left(\frac{n(x)}{n(x)-16}\right)}\)

OpenStudy (baru):

so we have \(\color{blue}{\displaystyle dy/dx= \left(\frac{n(x)}{n(x)-16}\right)}\) and \[\color{blue}{\displaystyle dy/dx=\left(\frac{n'(x)(n(x)-16) - (n'(x)n(x))}{(n(x)-16)^2}\right)x+\left(\frac{n(x)}{n(x)-16}\right)+n'(x) }\] therefore \[\color{blue}{\displaystyle \frac{n(x)}{n(x)-16} =\left(\frac{n'(x)(n(x)-16) - (n'(x)n(x))}{(n(x)-16)^2}\right)x+\left(\frac{n(x)}{n(x)-16}\right)+n'(x) }\]

OpenStudy (baru):

now luckily, the above expression simplifies nicely, and the n' terms cancel leaving \[16x = (16-n(x))^2\]

OpenStudy (baru):

so now we we isolate n(x) \[\pm 4\sqrt{x} = 16 - n(x) \\ n(x) =16 \pm 4\sqrt{x}\]

OpenStudy (baru):

substitute this n(x) in our eariler expression \(\color{blue}{\displaystyle dy/dx= \left(\frac{n(x)}{n(x)-16}\right)}\)

OpenStudy (baru):

we get \[\frac{dy}{dx} = \frac{4}{\sqrt{x}} +1 \] or \[\frac{dy}{dx} = -\frac{4}{\sqrt{x}} +1 \]

OpenStudy (baru):

in the graph you plotted, all the lines have a negative slope, and the first equation gives only positive values, so we will discard it

OpenStudy (baru):

\(\frac{dy}{dx} = -\frac{4}{\sqrt{x}} +1\) integrate both sides to get \[y=-8 \sqrt x+x+c\]

OpenStudy (baru):

now all thats left is to find c

OpenStudy (baru):

to find c, we need to find a point that definitely lies on the curve, so that we can substiute x and y to find c

OpenStudy (baru):

use \(n=16-4\sqrt{x}\) let n =8 the we get x=4 now substiute n=8 and x=4 in \(\color{blue}{\displaystyle y=\left(\frac{n(x)}{n(x)-16}\right)x+n (x) }\) to get y=4 so x=4 and y=4 is a point on y=f(x) substitute these values to find c. (you should get c=16)

OpenStudy (baru):

thats it :D that was incredibly fun to solve... and i was overjoyed when i plotted it found that it actually fits!! i really hope that its correct XD thanks @idku, this made my day.